A report in a research journal states that the average weight loss of people on a certain drug is 18 lbs with a margin of error of \pm5 lbs with confidence level C = 95%.

(a) According to this information, the mean weight loss of people on this drug, \mu , could be as low as________ lbs.

(b) If the study is repeated, how large should the sample size be so that the margin of error would be less than 2.5 lbs? (Assume \sigma = \ 8 lbs.)
ANSWER: ______
Answer #1

Solution:

(a) 18 - 5 = 13
as low as 13 lbs.
According to this information, the mean weight loss of people on this drug, \mu , could be as low as 13 lbs.
----------------------------------
(b) Margin of error = 2.5
=> z *σ/√n = 2.5
=> 1.96 * 8/√n = 2.5
=> n = (1.96 * 8/2.5)2
=> n = 40